### v1

# Refinement-based Christoffel sampling for least squares approximation in non-orthogonal bases

Abstract
We introduce a refinement-based Christoffel sampling (RCS) algorithm for least squares approximation in the span of a given, generally non-orthogonal set of functions $\Phi_{n}=\{\phi_{1},\dots,\phi_{n}\}$. A standard sampling strategy for this problem is Christoffel sampling, which achieves near-best approximations in probability using only $\mathcal{O}(n\log(n))$ samples. However, it requires i.i.d. sampling from a distribution whose density is proportional to the inverse Christoffel function $k_{n}$, the computation of which requires an orthonormal basis. As a result, existing approaches for non-orthogonal bases $\Phi_{n}$ typically rely on costly discrete orthogonalization. We propose a new iterative algorithm, inspired by recent advances in approximate leverage score sampling, that avoids this bottleneck. Crucially, while the computational cost of discrete orthogonalization grows proportionally with $\|k_{n}\|_{L^{\infty}(X)}$, the cost of our approach increases only logarithmically in $\|k_{n}\|_{L^{\infty}(X)}$. In addition, we account for finite-precision effects by considering a numerical variant of the Christoffel function, ensuring that the algorithm relies only on computable quantities. Alongside a convergence proof, we present extensive numerical experiments demonstrating the efficiency and robustness of the proposed method.
keywords:
Weighted least squares, sampling strategies, Christoffel function, non-orthogonal expansions, leverage scores
{MSCcodes}
41A10, 41A63, 42C15, 65C05
1 Introduction
We consider the problem of approximating a function $f\in L^{2}_{\rho}(X)$, defined on $X\subseteq\mathbb{R}^{d}$, using weighted least squares fitting in the span of a given basis $\Phi_{n}=\{\phi_{1},\dots,\phi_{n}\}$
(1) $\tilde{f}\in\operatorname*{arg\,min}_{v\in\operatorname{span}(\Phi_{n})}\sum_{k=1}^{m}w_{k}\left|f(x_{k})-v(x_{k})\right|^{2},$
where the set of sample points $\{x_{k}\}_{k=1}^{m}$ and the weights $\{w_{k}\}_{k=1}^{m}$ can be chosen freely. The objective is to attain near-best approximation errors while using a small number of sample points. For simplicity of exposition, we assume that the functions $\phi_{i}$ are linearly independent so that the dimension of $\operatorname{span}(\Phi_{n})$ equals $n$; however, the discussed methods also apply to linearly dependent sets of functions.
A common strategy is to use sample points that are drawn i.i.d. from some sampling distribution, allowing for an analysis using tools from random matrix theory. This led to the sampling distribution $\mu_{n}$, first proposed for (total degree) polynomial spaces [21] and later generalized [9], which achieves a near-optimal approximation error in probability using only $m=\mathcal{O}(n\log n)$ samples. This is close to the optimal case of interpolation, which corresponds to $m=n$. The sampling distribution $\mu_{n}$ is defined by
$w_{n}d\mu_{n}=d\rho\qquad\text{with}\qquad w_{n}=\frac{n}{\sum_{i=1}^{n}\lvert u_{i}\rvert^{2}},$
where $\{u_{i}\}_{i=1}^{n}$ is an orthonormal basis for $\operatorname{span}(\Phi_{n})$. The function
$k_{n}=\sum_{i=1}^{n}\lvert u_{i}\rvert^{2}$
is known as the inverse of the Christoffel function [31] and, hence, this sampling strategy is often referred to as Christoffel sampling [1].
An important practical challenge is that sampling from $\mu_{n}$ requires evaluating the Christoffel function, which in turn typically requires access to an orthonormal basis. We are particularly interested in settings where such a basis is not available, i.e., when the given basis $\Phi_{n}$ is non-orthogonal. Non-orthogonal bases often arise when one incorporates prior knowledge about the target functions, such as known singularities, oscillations or other features that a standard polynomial basis would fail to capture efficiently. They also naturally appear when working on irregular domains. In addition, non-orthogonal bases can result from adaptive or problem-driven constructions, such as bases learned using machine learning techniques.
In earlier work [15, 2, 30], this issue was addressed by discretely orthogonalizing the basis $\Phi_{n}$ on a dense grid. More specifically, we can equivalently characterize the inverse Christoffel function $k_{n}$ using the non-orthogonal basis $\Phi_{n}$ by
(2) $k_{n}(x)=\phi(x)^{*}G^{-1}\phi(x),$
where $\phi(x)=\begin{bmatrix}\phi_{1}(x)&\dots&\phi_{n}(x)\end{bmatrix}^{\top}$ and $G$ is the continuous Gram matrix associated with $\Phi_{n}$, defined by
$(G)_{ij}=\langle\phi_{i},\phi_{j}\rangle_{L^{2}_{\rho}(X)}.$
Discrete orthogonalization amounts to replacing $G$ by a discrete Gram matrix $G_{\ell}$, such as
$(G_{\ell})_{ij}=\sum_{k=1}^{\ell}\frac{1}{\ell}\phi_{i}(t_{k})\overline{\phi_{j}(t_{k})},$
where the points $\{t_{k}\}_{k=1}^{\ell}$ are sampled i.i.d. from the measure $\rho$. Following [15, 1], one needs $\ell=\mathcal{O}(\|k_{n}\|_{L^{\infty}(X)}\log(n))$ for $G_{\ell}$ to approximate $G$ well with high probability. Although this is typically an offline cost—computed once for every $\Phi_{n}$ and independently of the function to be approximated—it can still be prohibitively expensive, especially when the Christoffel function $k_{n}$ is sharply peaked. Notably, this is precisely the regime where Christoffel sampling offers significant advantages over standard least squares fitting using samples drawn from $\rho$. When $\ell$ is too small, the efficiency of the resulting Christoffel sampling procedure is limited.
We adopt a different strategy inspired by [10]; Section 1.2 explains how this discrete method relates to our continuous setting. We aim to construct a function $u$ that satisfies $u(x)\geq k_{n}(x),\forall x\in X$ while $\|u\|_{L^{1}_{\rho}(X)}\leq Cn$ for some small constant $C\geq 1$. When such a function is known, it suffices to draw $m=\mathcal{O}(\|u\|_{L^{1}_{\rho}(X)}\log(n))=\mathcal{O}(n\log(n))$ samples from the sampling distribution $\mu_{u}$ defined by
$w_{u}d\mu_{u}=d\rho\qquad\text{with}\qquad w_{u}=\frac{\|u\|_{L^{1}_{\rho}(X)}}{u}$
in order to obtain near-optimal weighted least squares approximations in expectation. We construct such a function $u$ efficiently using an iterative method in which, at each step, we sample points from the current $u$ and thereafter use these samples to update and improve $u$.
In particular, we show that $\|u\|_{L^{1}_{\rho}(X)}$ decreases by a constant factor in each iteration until it is close to $\|k_{n}\|_{L^{1}_{\rho}(X)}=n$, using only $\mathcal{O}(n\log(n))$ samples per step. Starting from the initialization $u(x)=\|k_{n}\|_{L^{\infty}(X)}$, it follows that the total number of samples required to obtain an adequate $u$ is of the order $\mathcal{O}(\log(\|k_{n}\|_{L^{\infty}(X)})n\log(n))$. Therefore, in contrast to the method described above, the computational complexity of this approach depends only logarithmically on $\|k_{n}\|_{L^{\infty}(X)}$, rather than linearly. Inspired by [10, Algorithm 3], the algorithm is termed refinement-based Christoffel sampling (RCS).
1.1 Effects of finite-precision arithmetic
Another issue related to least squares fitting with non-orthogonal bases is the influence of numerical rounding errors. Even if a basis is linearly independent from an analytical perspective, it can be indistinguishable from a linearly dependent set in finite-precision arithmetic. Numerically computed least squares fits using such a basis behave either explicitly or implicitly as $\ell^{2}$-regularized approximations with a regularization parameter $\epsilon\propto\epsilon_{\text{mach}}\sqrt{\|G\|_{2}}$ due to rounding errors [4, 5, 23]. Whereas this numerical regularization has a negligible influence if the basis is (nearly) orthogonal, it becomes significant when $\kappa(G)\geq 1/\epsilon_{\text{mach}}^{2}$, in which case $\Phi_{n}$ is called numerically redundant [23].
Using similar tools from random matrix theory, it was found in [23] that the sampling distribution $\mu_{n}^{\epsilon}$ defined by
$w_{n}^{\epsilon}d\mu_{n}^{\epsilon}=d\rho\qquad\text{with}\qquad w_{n}^{\epsilon}=\frac{n^{\epsilon}}{k_{n}^{\epsilon}}$
is of interest in this case. Here,
$k_{n}^{\epsilon}(x)=\phi(x)^{*}(G+\epsilon^{2}I)^{-1}\phi(x)\qquad\text{and}\qquad n^{\epsilon}=\sum_{i=1}^{n}\frac{\lambda_{i}(G)}{\lambda_{i}(G)+\epsilon^{2}}=\|k_{n}^{\epsilon}\|_{L^{1}_{\rho}(X)}$
are called the numerical inverse Christoffel function and the numerical dimension, respectively, and $\lambda_{i}(G)$ denotes the $i$-th eigenvalue of $G$. The numerical dimension $n^{\epsilon}$ now plays the role of the original dimension $n$, so that only $\mathcal{O}(n^{\epsilon}\log(n^{\epsilon}))$ samples are required for accurate least squares fitting, compared to $\mathcal{O}(n\log(n))$ when rounding errors are not accounted for. Note that $n^{\epsilon}\leq n$, reflecting the principle that numerical rounding errors can only reduce the dimension of the approximation space.
Even though in many practical scenarios one has $n^{\epsilon}\propto n$, such that accounting for the influence of rounding errors does not have a significant impact on the sampling complexity, the algorithm is based on the numerical inverse Christoffel function. A few observations clarify why this choice is necessary. First, it ensures that our algorithm only works with computable quantities. By constrast, evaluating the standard Christoffel function (2) may require inverting an arbitrarily ill-conditioned matrix, which is not feasible in finite precision. Moreover, observe that
$G\preceq G+\epsilon^{2}I\preceq\left(1+\frac{\epsilon^{2}}{\lambda_{n}(G)}\right)G\quad\Rightarrow\quad\left(1+\frac{\epsilon^{2}}{\lambda_{n}(G)}\right)^{-1}k_{n}\leq k_{n}^{\epsilon}\leq k_{n},$
and, similarly for the numerical dimension $n^{\epsilon}$,
$\frac{\lambda_{n}(G)}{\lambda_{n}(G)+\epsilon^{2}}\leq\frac{\lambda_{i}(G)}{\lambda_{i}(G)+\epsilon^{2}}\leq 1\quad\Rightarrow\quad\left(1+\frac{\epsilon^{2}}{\lambda_{n}(G)}\right)^{-1}n\leq n^{\epsilon}\leq n.$
Thus, $k_{n}^{\epsilon}$ and $n^{\epsilon}$ deviate from $k_{n}$ and $n$, respectively, by more than a factor of two only when $\lambda_{n}(G)\leq\epsilon^{2}\propto\lambda_{1}(G)\epsilon_{\text{mach}}^{2}$, i.e., when $\Phi_{n}$ is (nearly) numerically redundant. Conversely, when $\Phi_{n}$ is close to orthogonal, they behave almost identically to $k_{n}$ and $n$. This shows that the algorithm remains meaningful even when rounding errors are negligible. Finally, the numerical regularization parameter $\epsilon$ plays a crucial role in the convergence proof of the algorithm: the proof of Lemma 4.3 heavily relies on the assumption that $\epsilon>0$.
1.2 Connection to the discrete setting
Christoffel sampling is widely recognized as being closely related to leverage score sampling, a well-established technique in numerical linear algebra [18, 1]. Given an $m\times n$ matrix $A$ with $m>n$, the $j$-th statistical leverage score $\ell_{j}$ is defined by
$\ell_{j}=\sum_{i=1}^{n}\lvert(U)_{ji}\rvert^{2}\qquad 1\leq j\leq m,$
where $U$ denotes the $m\times n$ matrix containing the left singular vectors of the matrix $A$ [16]. Leverage scores identify which rows of a matrix are most important and are typically used to subsample a tall-and-skinny matrix $A$ before solving a least squares problem $Ax\approx b$. Other applications include, but are not limited to, low-rank approximation, matrix completion and kernel-based machine learning [17].
From the definition of $\ell_{j}$, it is clear that leverage scores are a discrete analogue of the inverse Christoffel function $k_{n}$. In particular, the continuous variable $x\in X$ is substituted by a discrete index $j$, where $1\leq j\leq m$. Furthermore, the Christoffel function serves a similar role in approximation theory as the leverage scores in numerical linear algebra, namely in subsampling the $\infty$-dimensional regression problem
$\begin{bmatrix}\phi_{1}&\phi_{2}&\dots&\phi_{n}\end{bmatrix}c\approx f.$
The numerical Christoffel function $k_{n}^{\epsilon}$ can similarly be seen as a continuous analogue of ridge leverage scores [23], which arise in the context of $\ell^{2}$-regularized least squares problems [6].
Challenges similar to those encountered in Christoffel sampling also appear in the context of leverage scores. In particular, computing leverage scores requires the full singular value decomposition of $A$, which is as computationally expensive as solving the original least squares problem. Consequently, there has been substantial research on fast computation—or rather, approximation—of leverage scores. One approach introduced in [10, 11] consists of computing a tight upper bound to the (rigde) leverage scores via an efficient iterative scheme. This approach can be extended to the continuous setting and forms the foundation of this paper.
1.3 Other related work
Other methods have been proposed for efficiently computing the Christoffel function from a non-orthogonal basis. For example, in  [15, §4] a multilevel strategy is introduced that approximates the inverse Christoffel function through successive refinements based on nested subspaces. However, this approach relies on rather strong assumptions about these subspaces and does not account for the effects of finite-precision arithmetic. Another iterative method is presented in [33], which requires the strong assumption $\lambda_{\min}(G)\gg\epsilon_{\text{mach}}$—where $\lambda_{\min}$ indicates the smallest nonzero eigenvalue—for computability.
Moreover, we note that optimizations have been studied to improve upon the Christoffel sampling scheme. These strategies reduce the number of required samples from $\mathcal{O}(n\log(n))$ to $\mathcal{O}(n)$, typically by leaving the i.i.d. setting. For instance, in [7] a subsampling scheme is introduced that is able to compress a good sampling set, while the algorithm in [14] is designed to directly sample from a distribution that incorporates both the Christoffel function and effective resistance. Implementing the latter can be done via rejection sampling with draws from the Christoffel function, as mentioned in [14, §5].
The link between leverage scores and the Christoffel function is not new; a more extensive comparison can be found in [29, §2]. Furthermore, in the discrete setting many more algorithms have been proposed to efficiently approximate leverage scores, such as [16] which uses random sketching to speed up the computations. Viewing the object $\begin{bmatrix}\phi_{1}&\phi_{2}&\dots&\phi_{n}\end{bmatrix}$ as the limit of a tall-and-skinny matrix was popularized in the Matlab package chebfun, where it is referred to as a quasimatrix [19].
1.4 Overview of the paper
The remainder of the paper is organized as follows. In Section 2, the symbols and notation used throughout the paper are summarized. In Section 3, we introduce the proposed iterative algorithm, outline its main objective and provide an initial discussion of its computational performance. Section 4 presents a theoretical analysis of the method, including a proof of its convergence properties. Finally, Section 5 is devoted to extensive numerical experiments that illustrate the efficiency and robustness of our approach.
2 Symbols and notation
Here, we collect the main symbols used throughout the paper. Let $X\subseteq\mathbb{R}^{d}$ denote the domain of interest, equipped with a probability measure $\rho$. The inner product on $L^{2}_{\rho}(X)$ is defined by
$\langle f,g\rangle_{L^{2}_{\rho}(X)}=\int_{X}f\;\overline{g}\;d\rho.$
We introduce the following sampling measure based on a strictly positive function $u$:
(3) $w_{u}d\mu_{u}=d\rho\qquad\text{with}\qquad w_{u}=\frac{\|u\|_{L^{1}_{\rho}(X)}}{u}.$
Note that this corresponds to the choices
$x_{k}\overset{\mathrm{iid}}{\sim}\mu_{u}\qquad\text{and}\qquad w_{k}=w_{u}(x_{k})/m$
in (1). The set $\Phi_{n}=\{\phi_{1},\dots,\phi_{n}\}\subset L^{2}_{\rho}(X)$ is a given set of (non-orthogonal) basis functions. The vector of basis evaluations at a point $x\in X$ is denoted by $\phi(x)=\begin{bmatrix}\phi_{1}(x)&\dots&\phi_{n}(x)\end{bmatrix}^{\top}\in\mathbb{C}^{n}$. The continuous Gram matrix associated with $\Phi_{n}$ is an $n\times n$ matrix $G$ defined by
$(G)_{ij}=\langle\phi_{i},\phi_{j}\rangle_{L^{2}_{\rho}(X)}.$
The numerical inverse Christoffel function equals
$k_{n}^{\epsilon}(x)=\phi(x)^{*}(G+\epsilon^{2}I)^{-1}\phi(x)$
where $\epsilon>0$ models finite-precision rounding and denotes the Hermitian conjugate. The numerical dimension is defined by
$n^{\epsilon}=\sum_{i=1}^{n}\frac{\lambda_{i}(G)}{\lambda_{i}(G)+\epsilon^{2}},$
where $\lambda_{i}(G)$} are the eigenvalues of $G$. For Hermitian matrices $A$ and $B$, we use $A\preceq B$ to denote the Loewner order, i.e., BAB-A is positive semidefinite.
3 Description of the algorithm
3.1 Main objective
The goal is to identify a small number of samples $\{x_{k}\}_{k=1}^{m}$ and weights $\{w_{k}\}_{k=1}^{m}$ so that the discrete least squares fit to a function $f$ in a given basis $\Phi_{n}$ has near-optimal approximation error. Our approach to obtain such samples is to compute a function $u$ that satisfies
(4) $u(x)\geq k_{n}^{\epsilon}(x),\;\forall x\in X\qquad\text{while}\qquad\|u\|_{L^{1}_{\rho}(X)}\leq Cn^{\epsilon}$
for some small constant $C\geq 1$ and numerical parameter $\epsilon\propto\epsilon_{\text{mach}}\sqrt{\|G\|_{2}}$. In this section, we verify that samples drawn from the associated measure $\mu_{u}$ (3) indeed result in stable and accurate least squares fits, both when $\Phi_{n}$ is numerically redundant and when numerical rounding errors are negligble, i.e., if $\Phi_{n}$ is not too far from being orthogonal.
Most importantly, if the sample points $\{x_{k}\}_{k=1}^{m}$ are drawn independently from $\mu_{u}$, then, with high probability, the discrete Gram matrix $G_{m}$ defined by
$(G_{m})_{ij}=\sum_{k=1}^{m}\frac{w_{u}(x_{k})}{m}\phi_{i}(x_{k})\overline{\phi_{j}(x_{k})},$
closely approximates the continuous Gram matrix $G$ using only $m=\mathcal{O}(n^{\epsilon}\log(n^{\epsilon}))$ samples. This is formalized in Lemma 3.1, which generalizes [23, Theorem 10].
Lemma 3.1.
Let $\Phi_{n}=\{\phi_{i}\}_{i=1}^{n}\subset L^{2}_{\rho}(X)$ be a basis, where $X\subseteq\mathbb{R}^{d}$, $\epsilon\geq 0$, and $u$ be an integrable function satisfying
$u(x)\geq k_{n}^{\epsilon}(x)\qquad\forall x\in X.$
Assuming that $k_{n}^{\epsilon}(x)>0,\forall x\in X$, let $\Delta>0$ and $\gamma<1$ and generate
$m\geq c_{\Delta}\log(16n^{\epsilon}/\gamma)\|u\|_{L^{1}_{\rho}(X)}$
samples $\{x_{k}\}_{k=1}^{m}$ i.i.d. from $\mu_{u}$ (3), where $c_{\Delta}=\frac{1+2\Delta/3}{\Delta^{2}/2}$. Then, assuming $\|G\|_{2}\geq\epsilon^{2}$, one has
(5) $(1-\Delta)(G+\epsilon^{2}I)\preceq G_{m}+\epsilon^{2}I\preceq(1+\Delta)(G+\epsilon^{2}I)$
with probability at least $1-\gamma$.
Proof 3.2.
This follows from a straightforward generalization of the proof of [23, Theorem 10] using a generic $\Delta$ instead of $\Delta=1/2$, and from
$\|w_{u}k^{\epsilon}_{n}\|_{L^{\infty}(X)}=\operatorname*{ess\,sup}_{x\sim\rho}\frac{k_{n}^{\epsilon}(x)\|u\|_{L^{1}_{\rho}(X)}}{u(x)}\leq\|u\|_{L^{1}_{\rho}(X)}.$
Remark 3.3.
Throughout this paper we will assume that $k_{n}^{\epsilon}(x)>0,\forall x\in X$. This is a standard assumption in Christoffel sampling schemes and is satisfied if at every point $x\in X$ at least one basis function is nonzero. This condition is mild and, for example, holds whenever the constant function is included in the basis. An alternative is to slightly modify the Christoffel density, as explained in [1, §6.1].
Using [23, Theorem 2], this immediately implies that numerically computed least squares fits are stable and accurate.
Corollary 3.4.
Under the same conditions as Lemma 3.1, given a function $f\in L^{2}_{\rho}(X)$ one has
$\left\lVert f-\sum_{i=1}^{n}\hat{c}_{i}\phi_{i}\right\rVert_{L^{2}_{\rho}(X)}\leq\left(1+\frac{1+\sqrt{2}}{\sqrt{1-\Delta}}\right)\inf_{c\in\mathbb{C}^{n}}\left\lVert f-\sum_{i=1}^{n}c_{i}\phi_{i}\right\rVert_{L^{\infty}(X)}+\epsilon\|c\|_{2}$
with probability at least $1-\gamma$, where
$\hat{c}\in\operatorname*{arg\,min}_{c\in\mathbb{C}^{n}}\sum_{k=1}^{m}\frac{w_{u}(x_{k})}{m}\left|f(x_{k})-\sum_{i=1}^{n}c_{i}\phi_{i}(x_{k})\right|^{2}+\epsilon^{2}\|c\|_{2}^{2}.$
At first glance these results seem tailored to the case where $\Phi_{n}$ is numerically redundant, yet it is important to note that this is not the case. As mentioned in the introduction, the influence of the numerical regularization parameter $\epsilon$ is only significant when $\Phi_{n}$ is numerically redundant. Hence, the results from Lemma 3.1 are also useful in the more standard setting where numerical rounding errors are negligible. In particular, (5) implies that
(6) $\left(1-\widetilde{\Delta}\right)G\preceq G_{m}\preceq\left(1+\widetilde{\Delta}\right)G\qquad\text{with}\qquad\widetilde{\Delta}=\Delta\left(1+\frac{\epsilon^{2}}{\lambda_{1}(G)}\right),$
where the term $\epsilon^{2}/\lambda_{1}(G)$ is negligble when $\Phi_{n}$ is close to orthogonal. From here, it follows that the discrete least squares error is near-optimal using $\mathcal{O}(n^{\epsilon}\log(n^{\epsilon}))=\mathcal{O}(n\log(n))$ samples, by arguments analogous to those used in standard results such as [9, Theorem 2.1].
Corollary 3.5.
Under the same conditions as Lemma 3.1, given a function $f\in L^{2}_{\rho}(X)$ one has
$\left\lVert f-\tilde{f}\right\rVert_{L^{2}_{\rho}(X)}\leq\left(1+\frac{1}{\sqrt{1-\widetilde{\Delta}}}\right)\inf_{v\in\operatorname{span}(\Phi_{n})}\left\lVert f-v\right\rVert_{L^{\infty}(X)}$
with probability at least $1-\gamma$, where $\widetilde{\Delta}$ as in (6) and
$\tilde{f}\in\operatorname*{arg\,min}_{v\in\operatorname{span}(\Phi_{n})}\sum_{k=1}^{m}\frac{w_{u}(x_{k})}{m}\left|f(x_{k})-v(x_{k})\right|^{2}.$
3.2 Iterative scheme
The core strategy for constructing a function $u$ that satisfies (4) is an adaptation of the (ridge) leverage score sampling algorithm introduced in [10, 11] to the continuous setting. Suppose we have a candidate function $u$ that upper bounds $k_{n}^{\epsilon}$ but with $\|u\|_{L^{1}_{\rho}(X)}\gg n^{\epsilon}$. Drawing fewer than $\mathcal{O}(\|u\|_{L^{1}_{\rho}(X)}\log(n^{\epsilon}))$ samples from the associated measure $\mu_{u}$ (3) is generally insufficient for accurate least squares fitting. However, these samples can still be used to improve uu—that is, to construct a new function $u^{(new)}$ that still upper bounds $k_{n}^{\epsilon}$ but has a reduced $L_{\rho}^{1}(X)$-norm. This procedure is applied iteratively, producing successively tighter upper bounds to $k_{n}^{\epsilon}$ until the requirement (4) is met.
In each iteration, we update the function $u$ as follows:
(7) $u^{(new)}(x)=\min\left(u(x),\;(1+\Delta)\phi(x)^{*}(A^{*}A+\epsilon^{2}I)^{-1}\phi(x)\right).$
Here, $A$ consists of $\lceil\alpha m+1\rceil$ rows of the $m\times n$ matrix $A_{m}$ defined by
$(A_{m})_{kj}=\sqrt{\frac{w_{u}(x_{k})}{m}}\overline{\phi_{j}(x_{k})}=\frac{\overline{\phi_{j}(x_{k})}}{\sqrt{C_{1}u(x_{k})\log(n^{\epsilon})}}\qquad\text{where}\qquad x_{k}\overset{\mathrm{iid}}{\sim}\mu_{u}$
with
$m=C_{1}\|u\|_{L^{1}_{\rho}(X)}\log(n^{\epsilon})$
for some parameters $\Delta,\alpha\in(0,1]$ and $C_{1}\geq 1$. Importantly, since each row of $A_{m}$ is independent, constructing $A$ only requires drawing $\lceil\alpha m+1\rceil$ samples. In other words: $A_{m}$ is never formed explicitly.
Figure 1 illustrates how this construction ensures that $u^{(new)}$ remains an upper bound to $k_{n}^{\epsilon}$ with high probability. According to Lemma 3.1, for sufficiently large $C_{1}$,
$A_{m}^{*}A_{m}+\epsilon^{2}I\preceq(1+\Delta)(G+\epsilon^{2}I)$
with high probability. Since $A$ is a submatrix of $A_{m}$, it follows that
$A^{*}A+\epsilon^{2}I\preceq A_{m}^{*}A_{m}+\epsilon^{2}I\preceq(1+\Delta)(G+\epsilon^{2}I)$
and, hence,
$(1+\Delta)\phi(x)^{*}(A^{*}A+\epsilon^{2}I)^{-1}\phi(x)\geq k_{n}^{\epsilon}(x)\qquad\forall x\in X$
with high probability. Assuming $u\geq k_{n}^{\epsilon}$ holds with high probability, this ensures that $u^{(new)}\geq k_{n}^{\epsilon}$ with high probability as well.
[…]
Figure 1: Schematic displaying the theoretical construction that motivates the definition of AA and, more specifically, why $u^{(new)}\geq k_{n}^{\epsilon}$ with high probability. Importantly, the matrix $A_{m}$ is never formed explicitly, so that only $\lceil\alpha m+1\rceil$ samples are drawn in each step.
How do we choose $\alpha$ so that $u$ improves significantly in each iteration while keeping the cost moderate? This is investigated in Section 4 and, more specifically, in Theorem 4.1, which provides a bound on $\|u^{(new)}\|_{L^{1}_{\rho}(X)}$ as a function of $\alpha$. If we choose
$\alpha=\frac{C_{2}n^{\epsilon}}{C_{1}\|u\|_{L^{1}_{\rho}(X)}}$
for some constant $C_{2}\geq 1$, then the amount of samples drawn in each iteration equals
$\lceil\alpha m+1\rceil=\lceil C_{2}n^{\epsilon}\log(n^{\epsilon})+1\rceil=\mathcal{O}(n^{\epsilon}\log(n^{\epsilon}))$
while
$\|u^{(new)}\|_{L^{1}_{\rho}(X)}\leq\left(\frac{2}{1-\Delta}\right)\frac{n^{\epsilon}}{\alpha}=\left(\frac{2}{1-\Delta}\right)\frac{C_{1}}{C_{2}}\|u\|_{L^{1}_{\rho}(X)}$
with high probability. Thus, for given $C_{1}$ and $\Delta$, choosing $C_{2}$ appropriately ensures a geometric decrease in $L_{\rho}^{1}(X)$-norm of $u$, using only $\mathcal{O}(n^{\epsilon}\log(n^{\epsilon}))$ samples in each iteration. Note that we don’t have to compute $m$ or $\alpha$; we simply draw $\lceil\alpha m+1\rceil$ samples in each iteration.
The complete procedure is given in Algorithm 1. A couple of remarks are in order. First, the bound above only applies when $\lceil\alpha m+1\rceil\leq m$, which roughly translates to $\alpha\leq 1$. Hence, we terminate the algorithm when $\alpha>1$, i.e., when
$\|u\|_{L^{1}_{\rho}(X)}<\frac{C_{2}n^{\epsilon}}{C_{1}}.$
At this point, the requirements in (4) are satisfied for $C=C_{2}/C_{1}$ with high probability, implying that the number of samples outputted by the algorithm can be bounded from above by
$C_{1}\|u\|_{L^{1}_{\rho}(X)}\log(n^{\epsilon})\leq C_{2}n^{\epsilon}\log(n^{\epsilon}).$
Second, we exploit that $u$ can be evaluated efficiently for multiple values of $x$ via
(8) $\phi(x)^{*}(A^{*}A+\epsilon^{2}I)^{-1}\phi(x)=\left\lVert(R^{*})^{-1}\phi(x)\right\rVert_{2}^{2},$
where
$QR=\begin{bmatrix}A\\ \epsilon I_{n\times n}\end{bmatrix}$
is a (thin) QR-decomposition. Hence, $u$ is updated by storing $R$.
Algorithm 1 Refinement-based Christoffel Sampling (RCS)
$\Phi_{n}\subset L^{2}_{\rho}(X)$, (upper bounds on) $\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}$ and $n^{\epsilon}$
$u(x)\leftarrow\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}$
while $\|u\|_{L^{1}_{\rho}(X)}>\tfrac{C_{2}}{C_{1}}n^{\epsilon}$ do
  Draw $\lceil C_{2}n^{\epsilon}\log(n^{\epsilon})+1\rceil$ samples $x_{k}$ i.i.d. from $\mu_{u}$ (3)
  Construct matrix AA: $(A)_{kj}\leftarrow\overline{\phi_{j}(x_{k})}/\sqrt{C_{1}u(x_{k})\log(n^{\epsilon})}$
Update $u$ (7) via (8) by computing the thin QR-decomposition of $\begin{bmatrix}A\\ \epsilon I_{n\times n}\end{bmatrix}$
  Estimate $\|u\|_{L^{1}_{\rho}(X)}$
end while
Draw $C_{1}\|u\|_{L^{1}_{\rho}(X)}\log(n^{\epsilon})$ samples $x_{k}$ i.i.d. from $\mu_{u}$ (3)
Compute the weights $w_{k}=1/(C_{1}u(x_{k})\log(n^{\epsilon}))$
return sample points $\{x_{1},x_{2},\dots\}$ and weights $\{w_{1},w_{2},\dots\}$
3.3 Discussion
The computational cost of the algorithm is
$\underbrace{\mathcal{O}(\log(\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}))}_{\text{\# iterations}}\;\times\underbrace{\mathcal{O}(n^{\epsilon}\log(n^{\epsilon}))}_{\text{\# samples per iteration}}\times\;\underbrace{\mathcal{O}(n^{2}\log(\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}))}_{\text{cost of evaluating $u$}}\;\;\text{flops}$
$=\;$ $\mathcal{O}(n^{3}\log(n)\log(\|k_{n}^{\epsilon}\|_{L^{\infty}(X)})^{2})\;\;\text{flops}$
where we used $n^{\epsilon}\leq n$. Here, the number of iterations depends logarithmically on $\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}$, since $\|u\|_{L^{1}_{\rho}(X)}$ decreases geometrically and equals $\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}$ at initialization. We also assume that drawing one sample from $\mu_{u}$ (3) requires a constant number of evaluations of the function $u$, as discussed further below.
One evaluation of $u$ in iteration $i$ corresponds to computing
$u(x)=\min\left(\|k_{n}^{\epsilon}\|_{L^{\infty}(X)},\;\left\lVert\left(R_{(1)}^{*}\right)^{-1}\phi(x)\right\rVert_{2}^{2},\;\dots,\;\left\lVert\left(R_{(i-1)}^{*}\right)^{-1}\phi(x)\right\rVert_{2}^{2}\right),$
where each system solve requires $\mathcal{O}(n^{2})$ flops since $R$ is upper-triangular. We find that the computational cost of evaluating $u$ can be reduced in practice, as discussed in Section 5. The total cost of computing the $R$-factors equals
$\underbrace{\mathcal{O}(\log(\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}))}_{\text{\# iterations}}\;\times\underbrace{\mathcal{O}(n^{\epsilon}\log(n^{\epsilon})n^{2})}_{\text{\parbox{71.13188pt}{\centering cost of QR \\ \vskip-2.84526ptdecomposition\@add@centering}}}=\mathcal{O}(n^{3}\log(n)\log(\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}))\;\text{flops},$
as this only needs to be done once per iteration.
We now discuss the assumptions underlying the algorithm. A further discussion on the practical implementation of the algorithm can be found in Section 5.
Known upper bound on $\boldsymbol{\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}}$
For polynomial spaces, the growth of the Christoffel function on irregular domains has been studied extensively [27, 15]. Even if one does not have a tight upper bound, overestimating it has little impact since the computational complexity depends only logarithmically on this bound.
Known upper bound on $\boldsymbol{n^{\epsilon}}$
A natural choice for an upper bound on the numerical dimension $n^{\epsilon}$ is $n$, since $n^{\epsilon}\leq n$ for all $\epsilon\geq 0$. In this case, the algorithm produces a sample set of size $\mathcal{O}(n\log(n))$, which suffices in most practical scenarios. If the basis $\Phi_{n}$ is highly numerically redundant, however, it can be advantageous to use a numerical dimension $n^{\epsilon}\ll n$.
If such an estimate is unknown, one might attempt to first run Algorithm 1 with $n$ as an upper bound, then estimate the numerical dimension using the resulting discrete Gram matrix $G_{m}$ and, if the estimate is significantly smaller than $n$, re-run the algorithm using this refined estimate. Unfortunately, this approach is unreliable. To see why, note that the discrete Gram $G_{m}$ is guaranteed to satisfy (5) with high probability. In the specific case where $G_{m}+\epsilon^{2}I=(1-\Delta)(G+\epsilon^{2}I)$, straightforward calculations show that
$\sum_{i=1}^{n}\frac{\lambda_{i}(G_{m})}{\lambda_{i}(G_{m})+\epsilon^{2}}=\operatorname{Tr}(G_{m}(G_{m}+\epsilon^{2}I)^{-1})=n^{\epsilon}-\frac{\Delta}{1-\Delta}\sum_{j=1}^{n}\frac{\epsilon^{2}}{\epsilon^{2}+\lambda_{i}(G)}.$
For highly redundant bases, this results in a significant underestimation of $n^{\epsilon}$, causing the algorithm to fail. Reliable estimates can instead be obtained when a condition analogous to (5) holds for some $\widetilde{\epsilon}\ll\epsilon$, i.e., by using multiple regularization parameters. We do not pursue this approach further.
Sampling from $\boldsymbol{\mu_{u}}$ and estimating $\boldsymbol{\|u\|_{L^{1}_{\rho}(X)}}$
The algorithm assumes that samples of $\mu_{u}$ and estimates of $\|u\|_{L^{1}_{\rho}(X)}$ can be obtained efficiently from evaluations of $u$. In practice, we find that Markov chain samplers yield satisfactory results, even though the samples are not independent. Moreover, these samplers can often run without knowledge of the normalization factor of $\mu_{u}$, i.e., $\|u\|_{L^{1}_{\rho}(X)}$, meaning that only a rough estimate of $\|u\|_{L^{1}_{\rho}(X)}$ is needed. This estimate is then only used to terminate Algorithm 1. We find that a simple Monte Carlo estimator or Matlab’s built-in integral functions work well.
4 Theoretical results
The algorithm presented in this paper is a continuous analogue of the iterative leverage score sampling algorithm proposed in [10, Algorithm 3], while incorporating $\ell^{2}$-regularization as in [11]. We heavily rely on this connection to prove the convergence behaviour of our algorithm.
The main goal is to prove that in each step of the algorithm, the updated function $u^{(new)}$ (7) remains an upper bound to $k_{n}^{\epsilon}$ while the $L^{1}_{\rho}(X)$-norm gets reduced compared to $u$. This is quantified in Theorem 4.1. For clarity, we are more explicit about the basis in our notation throughout this section. More specifically, we use
$k^{\epsilon}(x;\Psi)$ $=\psi(x)^{*}(G+\epsilon^{2}I)^{-1}\psi(x)$
$k^{\epsilon}(x;\Psi,M)$ $=\psi(x)^{*}(M+\epsilon^{2}I)^{-1}\psi(x)$
$n^{\epsilon}(\Psi)$ $=\int_{X}k^{\epsilon}(x;\Psi)d\rho=\sum_{i=1}^{n}\frac{\lambda_{i}(G)}{\lambda_{i}(G)+\epsilon^{2}}$
where $\psi(x)=\begin{bmatrix}\psi_{1}(x)&\dots&\psi_{n}(x)\end{bmatrix}^{\top}$ and $G$ is the continuous Gram matrix associated with $\Psi=\{\psi_{i}\}_{i=1}^{n}$.
Theorem 4.1.
Let $X\subseteq\mathbb{R}^{d}$ be compact, $\Phi=\{\phi_{i}\}_{i=1}^{n}\subset L^{2}_{\rho}(X)$ be a set of piecewise Lipschitz continuous functions, $\epsilon>0$ and assume $\inf_{x\in X}k^{\epsilon}(x;\Phi)>0$. Let $u$ be an integrable, piecewise Lipschitz function satisfying
$u(x)\geq k^{\epsilon}(x;\Phi),\quad\forall x\in X.$
For some $\alpha>0$, $\Delta>0$, $\gamma<1$, let
$m\geq c_{\Delta}\log\big(16n^{\epsilon}(\Phi)/\gamma\big)\int_{X}u\,d\rho,\qquad c_{\Delta}=\frac{1+2\Delta/3}{\Delta^{2}/2},$
for small enough $\alpha$ such that $\lceil\alpha m+1\rceil<m$. Draw i.i.d. samples $\{x_{k}\}_{k=1}^{\lceil\alpha m+1\rceil}$ from $\mu_{u}$ (3) and define the $\lceil\alpha m+1\rceil\times n$ matrix $A$ via
$(A)_{kj}=\sqrt{\frac{w_{u}(x_{k})}{m}}\overline{\phi_{j}(x_{k})}.$
Then, with probability at least $1-2\gamma$, the updated function $u^{(new)}$ (7) satisfies
(9) $u^{(new)}(x)$ $\geq k^{\epsilon}(x;\Phi),\quad\forall x\in X,$
(10) $\int_{X}u^{(new)}\,d\rho$ $\leq\left(\frac{2}{1-\Delta}\right)\frac{n^{\epsilon}(\Phi)}{\alpha}.$
Proof 4.2.
1. Establishing (9)
The matrix $A$ can be viewed as a submatrix of an $m\times n$ matrix $A_{m}$, equivalently defined by
$(A_{m})_{ij}=\sqrt{\frac{w_{u}(x_{k})}{m}}\overline{\phi_{j}(x_{k})}.$
Due to Lemma 3.1, we have $\frac{1}{1+\Delta}(A_{m}^{*}A_{m}+\epsilon^{2}I)\preceq G+\epsilon^{2}I$ with probability at least $1-\gamma$. Furthermore,
$A_{m}^{*}A_{m}-A^{*}A=\sum_{k=\lceil\alpha m+1\rceil+1}^{m}\frac{w(x_{k})}{m}\phi(x_{k})\phi(x_{k})^{*}\succeq 0\qquad\Rightarrow\qquad A^{*}A\preceq A_{m}^{*}A_{m}$
and, hence, $\frac{1}{1+\Delta}(A^{*}A+\epsilon^{2}I)\preceq G+\epsilon^{2}I$ with probability $1-\gamma$ as well. It follows that
$(1+\Delta)k^{\epsilon}\left(x;\Phi,A^{*}A\right)\geq k^{\epsilon}\left(x;\Phi,G\right)=k^{\epsilon}(x;\Phi)$
and, using the definition of $u^{(new)}$ and the fact that $u\geq k^{\epsilon}(x;\Phi)$,
$u^{(new)}(x)\geq k^{\epsilon}(x;\Phi)\qquad\forall x\in X$
as required.
2. Establishing (10)
It follows from Lemma 4.3, presented below, that for every $\delta_{1},\delta_{2}>0$ there exists a function $v$ such that
$\alpha u(x)+\delta_{1}\geq k^{\epsilon}(x;v\Phi),\quad\forall x\in X\qquad\text{and}\qquad\int_{X}\alpha u\;1_{\{v(x)\neq 1\}}\;d\rho\leq n^{\epsilon}(\Phi)+\delta_{2}.$
Using this $v$, we can derive that
[…]
$\leq\int_{X}u\;1_{\{v(x)\neq 1\}}\;d\rho+\int_{X}(1+\Delta)k^{\epsilon}\left(x;\Phi,A^{*}A\right)\;1_{\{v(x)=1\}}\;d\rho$
$\leq\frac{n^{\epsilon}(\Phi)+\delta_{2}}{\alpha}+\int_{X}(1+\Delta)k^{\epsilon}\left(x;\Phi,A^{*}A\right)\;1_{\{v(x)=1\}}\;d\rho$
(11) $=\frac{n^{\epsilon}(\Phi)+\delta_{2}}{\alpha}+\int_{X}(1+\Delta)k^{\epsilon}\left(x;v\Phi,A^{*}A\right)\;1_{\{v(x)=1\}}\;d\rho,$
where $v\Phi=\{v\phi_{i}\}_{i=1}^{n}$.
We proceed by bounding $k^{\epsilon}\left(x;v\Phi,A^{*}A\right)$. To this end, note that $0\leq v(x)\leq 1$ implies $V^{*}V\preceq A^{*}A$, where
$(V)_{kj}=\sqrt{\frac{w_{u}(x_{k})}{m}}v(x_{k})\overline{\phi_{j}(x_{k})}=v(x_{k})(A)_{kj}.$
Furthermore, since $\alpha u(x)\geq k^{\epsilon}(x;v\Phi)+\delta_{1}$, we have
(12) $(1-\Delta)(G^{v}+\epsilon^{2}I)\preceq\frac{m}{\lceil\alpha m+1\rceil}V^{*}V+\epsilon^{2}I\qquad\text{ where }(G^{v})_{ij}=\langle v\phi_{i},v\phi_{j}\rangle_{L^{2}_{\rho}(X)}$
with probability at least $1-\gamma$ using a similar reasoning as in Lemma 3.1 for the basis $v\Phi$. More specifically, note that the samples used to construct $V$ are drawn i.i.d. from $\mu_{u}$. In order for (12) to hold, we need that the number of samples $\lceil\alpha m+1\rceil$ satisfies
(13) $\begin{split}\lceil\alpha m+1\rceil&\geq c_{\Delta}\log(16n^{\epsilon}(v\Phi)/\gamma)\operatorname*{ess\,sup}_{x\sim\rho}\left(\frac{k^{\epsilon}(x;v\Phi)}{\alpha u(x)}\int_{X}\alpha u\;d\rho\right)\\ &=\alpha c_{\Delta}\log(16n^{\epsilon}(v\Phi)/\gamma)\int_{X}ud\rho\operatorname*{ess\,sup}_{x\sim\rho}\left(\frac{k^{\epsilon}(x;v\Phi)}{\alpha u(x)}\right).\end{split}$
Due to lemma B.1, it suffices that
$\lceil\alpha m+1\rceil\geq\alpha c_{\Delta}\log(16n^{\epsilon}(\Phi)/\gamma)\int_{X}ud\rho\operatorname*{ess\,sup}_{x\sim\rho}\left(\frac{k^{\epsilon}(x;v\Phi)}{\alpha u(x)}\right).$
By construction we have
$m\geq c_{\Delta}\log(16n^{\epsilon}(\Phi)/\gamma)\int_{X}ud\rho,$
and, furthermore,
$\operatorname*{ess\,sup}_{x\sim\rho}\left(\frac{k^{\epsilon}(x;v\Phi)}{\alpha u(x)}\right)$ $\leq\operatorname*{ess\,sup}_{x\sim\rho}\left(\frac{k^{\epsilon}(x;v\Phi)}{\alpha u(x)+\delta_{1}}\right)\operatorname*{ess\,sup}_{x\sim\rho}\left(\frac{\alpha u(x)+\delta_{1}}{\alpha u(x)}\right)$
$\leq 1+\frac{\delta_{1}}{\operatorname*{ess\,inf}_{x\sim\rho}\alpha u(x)}.$
If we consider some $\delta_{1}>0$ satisfying
$\delta_{1}\leq\frac{\operatorname*{ess\,inf}_{x\sim\rho}\alpha u(x)}{\alpha c_{\Delta}\log(16n^{\epsilon}(\Phi)/\gamma)\int_{X}ud\rho}=\frac{\operatorname*{ess\,inf}_{x\sim\rho}u(x)}{c_{\Delta}\log(16n^{\epsilon}(\Phi)/\gamma)\int_{X}ud\rho},$
it follows that (13) and, hence, (12) is satisfied. Note that this is possible since $\operatorname*{ess\,inf}_{x\sim\rho}u(x)>0$ due to $\inf_{x\in X}k^{\epsilon}(x;\Phi)>0$. In conclusion, we obtain the bound
$k^{\epsilon}\left(x;v\Phi,A^{*}A\right)$ $\leq k^{\epsilon}\left(x;v\Phi,V^{*}V\right)\qquad$ $(V^{*}V\preceq A^{*}A)$
$\leq\frac{m}{\lceil\alpha m+1\rceil}k^{\epsilon}\left(x;v\Phi,\frac{m}{\lceil\alpha m+1\rceil}V^{*}V\right)\qquad$ (Lemma B.3)
$\leq\frac{m}{\lceil\alpha m+1\rceil}\left(\frac{1}{1-\Delta}\right)k^{\epsilon}(x;v\Phi,G^{v})\qquad$ (12)
$\leq\frac{1}{\alpha}\left(\frac{1}{1-\Delta}\right)k^{\epsilon}(x;v\Phi).$
Using this bound in (11), we arrive at
$\int_{X}u^{(new)}\;d\rho$ $\leq\frac{n^{\epsilon}(\Phi)+\delta_{2}}{\alpha}+\int_{X}\frac{1}{\alpha}\left(\frac{1+\Delta}{1-\Delta}\right)k^{\epsilon}(x;v\Phi)\;1_{\{v(x)=1\}}\;d\rho$
$\leq\frac{n^{\epsilon}(\Phi)+\delta_{2}}{\alpha}+\frac{1}{\alpha}\left(\frac{1+\Delta}{1-\Delta}\right)n^{\epsilon}(v\Phi)$
$\leq\left(1+\left(\frac{1+\Delta}{1-\Delta}\right)\right)\frac{n^{\epsilon}(\Phi)}{\alpha}+\frac{\delta_{2}}{\alpha}$
$=\left(\frac{2}{1-\Delta}\right)\frac{n^{\epsilon}(\Phi)}{\alpha}+\frac{\delta_{2}}{\alpha},$
where in the third step we use Lemma B.1. Since this bound holds for every $\delta_{2}>0$, one has
$\int_{X}u^{(new)}\;d\rho$ $\leq\inf_{\delta_{2}>0}\left(\frac{2}{1-\Delta}\right)\frac{n^{\epsilon}(\Phi)}{\alpha}+\frac{\delta_{2}}{\alpha}=\left(\frac{2}{1-\Delta}\right)\frac{n^{\epsilon}(\Phi)}{\alpha}.$
3. Establishing the probability
[…]
$\mathbb{P}(E_{1}\cap E_{2})=1-\mathbb{P}(E_{1}^{c}\cup E_{2}^{c})\geq 1-\mathbb{P}(E_{1}^{c})-\mathbb{P}(E_{2}^{c})\geq 1-2\gamma.$
The proof of Theorem 4.1 relies heavily on the following lemma. This lemma asserts the existence of a weight function $v$ (with $0\leq v(x)\leq 1,\forall x\in X$) such that the numerical inverse Christoffel function $k_{n}^{\epsilon}$ associated with the weighted basis $v\Phi$ lies below a prescribed function $u$ (up to an arbitrarily small constant $\delta_{1}>0$). Moreover, the function $v$ does not distort the basis functions $\Phi$ too much, in the sense of (15). Note that without condition (15), condition (14) would be trivial to satisfy—for instance, by choosing $v=0$.
This lemma is closely related to discrete analogues, namely [10, Theorem 2] (for leverage scores) and [11, Lemma 6.2] (for ridge leverage scores with variable regularization parameter). On top of this connection, our proof proceeds by first discretizing the problem and then applying the proof techniques from [10, Theorem 2] and [11, Lemma 6.2]. Specifically, we require a discrete result on ridge leverage scores with a fixed regularization parameter, which is provided in Appendix A.
Lemma 4.3.
Let $X\subset\mathbb{R}^{n}$ be compact, $\{\phi_{i}\}_{i=1}^{n}\subset L^{2}_{\rho}(X)$ be a set of piecewise Lipschitz continuous functions, $u:X\to\mathbb{R}$ be a strictly positive and piecewise Lipschitz continuous function and $\epsilon>0$. Then, for every $\delta_{1},\delta_{2}>0$ there exists a function $v:X\to\mathbb{R}$ satisfying
$0\leq v(x)\leq 1\qquad\forall x\in X$
such that
(14) $k^{\epsilon}(x;v\Phi)\leq u(x)+\delta_{1}\qquad\forall x\in X$
where $v\Phi=\{v\phi_{i}\}_{i=1}^{n}$, and
(15) $\int_{X}u\;1_{\{v(x)\neq 1\}}\;d\rho\leq n^{\epsilon}(\Phi)+\delta_{2}.$
Proof 4.4.
We start by partitioning $X$ in $\ell$ $\rho$-measurable partitions $\{X_{k}\}_{k=1}^{\ell}$, identifying $\ell$ discretization points $\{x_{k}\}_{k=1}^{\ell}$ that satisfy $x_{k}\in X_{k}$, and constructing the matrix $B^{(\ell)}\in\mathbb{R}^{\ell\times n}$ via
$(B^{(\ell)})_{kj}=\sqrt{\rho(X_{k})}\kern 3.0pt\overline{\phi_{j}(x_{k})}.$
We assume that the partitions are chosen in such a way that both the functions $\{\phi_{i}\}_{i=1}^{n}$ and the function $u$ are continuous in all partitions $X_{k}$. Furthermore, we define
$h^{(\ell)}=\max_{1\leq k\leq\ell}\max_{x\in X_{k}}\left|x-x_{k}\right|,$
and
$u_{k}^{(\ell)}=\int_{X_{k}}u\;d\rho,\qquad 1\leq k\leq\ell.$
Since $u$ is strictly positive, $u_{k}^{(\ell)}>0$ for all $i$. From Lemma A.1, it follows that there exists a diagonal matrix $V^{(\ell)}\in\mathbb{R}^{\ell\times\ell}$ with $0\preceq V^{(\ell)}\preceq I$ such that
(16) $\tau_{k}^{\epsilon}(V^{(\ell)}B^{(\ell)})\leq u_{k}^{(\ell)},\qquad 1\leq k\leq\ell$
and
(17) $\sum_{V^{(\ell)}_{kk}\neq 1}u_{k}^{(\ell)}\leq n^{\epsilon}(B^{(\ell)}).$
Using this matrix $V^{(\ell)}$ we can define a piecewise continuous function $v^{(\ell)}$ by
$v^{(\ell)}(x)=\sum_{k=1}^{\ell}V^{(\ell)}_{kk}1_{X_{k}},$
where $1_{X_{k}}$ is the indicator function of $X_{k}$.
1. Establishing (15)
From (17), it follows immediately that
[…]
In order to relate $n^{\epsilon}(B^{(\ell)})$ to $n^{\epsilon}(\Phi)$, consider
(18) $\begin{split}\left|((B^{(\ell)})^{*}B^{(\ell)})_{ij}-(G)_{ij}\right|&=\left|\sum_{k=1}^{m}\rho(X_{k})\phi_{i}(x_{k})\overline{\phi_{j}(x_{k})}-\int_{X}\phi_{i}\overline{\phi_{j}}\;d\rho\right|\\ &\leq\sum_{k=1}^{m}\left|\rho(X_{k})\phi_{i}(x_{k})\overline{\phi_{j}(x_{k})}-\int_{X_{k}}\phi_{i}\overline{\phi_{j}}\;d\rho\right|\\ &=\sum_{k=1}^{m}\left|\int_{X_{k}}\phi_{i}(x_{k})\overline{\phi_{j}(x_{k})}-\phi_{i}\overline{\phi_{j}}\;d\rho\right|\\ &\leq C\rho(X)h^{(\ell)},\end{split}$
for some $C>0$, where in the last step we use that the functions $\phi_{i}$ are Lipschitz continuous in every partition and bounded. Using [25, §21-1], it follows that there exists a permutation $\tau$ over $\{1,2,\dots,n\}$ such that
$\lvert\lambda_{i}(G)-\lambda_{\tau(i)}((B^{(\ell)})^{*}B^{(\ell)})\rvert\leq\|G-(B^{(\ell)})^{*}B^{(\ell)}\|_{F}\leq Cn\rho(X)h^{(\ell)},\qquad 1\leq i\leq n$
and, hence,
$n^{\epsilon}(B^{(\ell)})\leq n^{\epsilon}(\Phi)+Cn^{2}\rho(X)h^{(\ell)}\epsilon^{-2}$
since $\lambda/(\lambda+\epsilon^{2})$ is monotonically increasing for nonnegative $\lambda$ and
$\frac{\lambda+\delta}{\lambda+\delta+\epsilon^{2}}\leq\frac{\lambda}{\lambda+\epsilon^{2}}+\epsilon^{-2}\delta$
for $\delta\geq 0$. Because it is possible to construct partitions such that $h^{(\ell)}\to 0$ as $\ell\to\infty$, we obtain that for every $\delta_{2}>0$ there exists a sufficiently large $\ell$ such that
$\int_{X}u\;1_{\{v(x)\neq 1\}}\;d\rho\leq n^{\epsilon}(\Phi)+\delta_{2}$
for $v=v^{(\ell)}$, as required.
2. Establishing (14)
We divide (16) by $\rho(X_{k})$ and consider
(19) $\begin{split}\tau^{\epsilon}_{i}(V^{(\ell)}B^{(\ell)})/\rho(X_{k})&\leq u_{k}^{(\ell)}/\rho(X_{k})\\ &\Updownarrow\\ k^{\epsilon}(x;v^{(\ell)}\Phi)+\underbrace{\tau^{\epsilon}_{k}(V^{(\ell)}B^{(\ell)})/\rho(X_{k})-k^{\epsilon}(x;v^{(\ell)}\Phi)}_{\text{error 1}}&\leq u(x)+\underbrace{u_{k}^{(\ell)}/\rho(X_{k})-u(x)}_{\text{error 2}}\end{split}$
for $x\in X_{k}$. We will show that both errors in the inequality above go to zero as $h^{(\ell)}$ goes to zero, for every $k$. Because it is possible to construct partitions such that $h^{(\ell)}\to 0$ as $\ell\to\infty$, we obtain that for every $\delta_{1}>0$ there exists a sufficiently large $\ell$ such that
$k^{\epsilon}(x;v\Phi)\leq u(x)+\delta_{1},\qquad\forall x\in X$
for $v=v^{(\ell)}$, as required
The first error can be bounded by splitting it as
(20) $\begin{split}\left|\tau^{\epsilon}_{k}(V^{(\ell)}B^{(\ell)})/\rho(X_{k})-k^{\epsilon}(x;v^{(\ell)}\Phi)\right|\;\leq\;&\left|\tau^{\epsilon}_{k}(V^{(\ell)}B^{(\ell)})/\rho(X_{k})-k\left(x;v^{(\ell)}\Phi,G^{v^{(\ell)}}_{\ell}\right)\right|\\ &+\left|k\left(x;v^{(\ell)}\Phi,G^{v^{(\ell)}}_{\ell}\right)-k^{\epsilon}(x;v^{(\ell)}\Phi)\right|\end{split}$
where $G^{v^{(\ell)}}_{\ell}=(V^{(\ell)}B^{(\ell)})^{*}V^{(\ell)}B^{(\ell)}$. Since $x\in X_{k}\Rightarrow v^{(\ell)}(x)=V^{(\ell)}_{kk}$, we can bound the first term by
$\left|\tau_{k}^{\epsilon}(V^{(\ell)}B^{(\ell)})/\rho(X_{k})-k\left(x;v^{(\ell)}\Phi,G^{v^{(\ell)}}_{\ell}\right)\right|$
$=\left|v^{(\ell)}(x_{k})^{2}\phi(x_{k})^{*}(G^{v^{(\ell)}}_{\ell}+\epsilon^{2}I)^{-1}\phi(x_{k})-v^{(\ell)}(x)^{2}\phi(x)^{*}(G^{v^{(\ell)}}_{\ell}+\epsilon^{2}I)^{-1}\phi(x)\right|$
$=\left(V^{(\ell)}_{kk}\right)^{2}\left|(\phi(x_{k})-\phi(x))^{*}(G^{v^{(\ell)}}_{\ell}+\epsilon^{2}I)^{-1}(\phi(x_{k})+\phi(x))\right|$
$\leq\|\phi(x_{k})-\phi(x)\|_{2}\|(G^{v^{(\ell)}}_{\ell}+\epsilon^{2}I)^{-1}\|_{2}\|\phi(x_{k})+\phi(x)\|_{2}$
$\leq Cn^{2}h^{(\ell)}\epsilon^{-2},$
for some $C>0$, where in the last step we use that the functions $\phi_{i}$ are Lipschitz continuous in $X_{k}$ and bounded. This term indeed goes to zero as $h^{(\ell)}\to 0$.
In order to bound the second part of (20), define $G^{v^{(\ell)}}\in\mathbb{C}^{n\times n}$ as
$\left(G^{v^{(\ell)}}\right)_{ij}=\langle v^{(\ell)}\phi_{i},v^{(\ell)}\phi_{j}\rangle_{L^{2}_{\rho}(X)}$
and consider
$\left|\left(G^{v^{(\ell)}}_{\ell}\right)_{ij}-\left(G^{v^{(\ell)}}\right)_{ij}\right|$ $=\left|\sum_{k=1}^{\ell}\rho(X_{k})v^{(\ell)}(x_{k})^{2}\phi_{i}(x_{k})\overline{\phi_{j}(x_{k})}-\int_{X}(v^{(\ell)})^{2}\phi_{i}\overline{\phi_{j}}\;d\rho\right|$
$\leq\sum_{k=1}^{\ell}\left|\rho(X_{k})v^{(\ell)}(x_{k})^{2}\phi_{i}(x_{k})\overline{\phi_{j}(x_{k})}-\int_{X_{k}}(v^{(\ell)})^{2}\phi_{i}\overline{\phi_{j}}\;d\rho\right|$
$=\sum_{k=1}^{\ell}\left(V^{(\ell)}_{kk}\right)^{2}\left|\rho(X_{k})\phi_{i}(x_{k})\overline{\phi_{j}(x_{k})}-\int_{X_{k}}\phi_{i}\overline{\phi_{j}}\;d\rho\right|$
$\leq\sum_{k=1}^{\ell}\left|\rho(X_{k})\phi_{i}(x_{k})\overline{\phi_{j}(x_{k})}-\int_{X_{k}}\phi_{i}\overline{\phi_{j}}\;d\rho\right|$
$\leq C\rho(X)h^{(\ell)},$
for some $C>0$, where the last step is equivalent to the steps in (18). This derivation indeed allows us to bound the second term of (20) for all $x\in X_{k}$, namely
$\left|k\left(x;v^{(\ell)}\Phi,G^{v^{(\ell)}}_{\ell}\right)-k^{\epsilon}(x;v^{(\ell)}\Phi)\right|$
$=\left|v^{(\ell)}(x)^{2}\phi(x)^{*}(G^{v^{(\ell)}}_{\ell}+\epsilon^{2}I)^{-1}\phi(x)-v^{(\ell)}(x)^{2}\phi(x)^{*}(G^{v^{(\ell)}}+\epsilon^{2}I)^{-1}\phi(x)\right|$
$=\left(V_{kk}^{(\ell)}\right)^{2}\left|\phi(x)^{*}(G^{v^{(\ell)}}_{\ell}+\epsilon^{2}I)^{-1}\phi(x)-\phi(x)^{*}(G^{v^{(\ell)}}+\epsilon^{2}I)^{-1}\phi(x)\right|$
$\leq\left|\phi(x)^{*}(G^{v^{(\ell)}}_{\ell}+\epsilon^{2}I)^{-1}\phi(x)-\phi(x)^{*}(G^{v^{(\ell)}}+\epsilon^{2}I)^{-1}\phi(x)\right|$
$=\left|\phi(x)^{*}((G^{v^{(\ell)}}_{\ell}+\epsilon^{2}I)^{-1}((G^{v^{(\ell)}}+\epsilon^{2}I)-(G^{v^{(\ell)}}_{\ell}+\epsilon^{2}I))(G^{v^{(\ell)}}+\epsilon^{2}I)^{-1})\phi(x)\right|$
$\leq\|\phi(x)\|^{2}_{2}\|(G^{v^{(\ell)}}_{\ell}+\epsilon^{2}I)^{-1}\|_{2}\|G^{v^{(\ell)}}-G^{v^{(\ell)}}_{\ell}\|_{2}\|(G^{v^{(\ell)}}+\epsilon^{2}I)^{-1}\|_{2}$
$\leq C\epsilon^{-4}n^{5/2}\rho(X)h^{(\ell)},$
for some $C>0$, where in the last step we use that the functions $\phi_{i}$ are bounded and $\|M\|_{2}\leq\sqrt{n}\max_{1\leq i\leq n}\sum_{j=1}^{n}\lvert(M)_{ij}\rvert$ for $M\in\mathbb{C}^{n\times n}$ [25, §50-4]. This term also goes to zero as $h^{(\ell)}\to 0$ and, hence, the first error in (19) goes to zero as $h^{(\ell)}\to 0$.
Similarly for the second error in (19), we get for all $x\in X_{k}$ that
$\left|u_{i}^{(\ell)}/\rho(X_{k})-u(x)\right|$ $=\left|\int_{X_{k}}u(y)-u(x)\;d\rho(y)\right|/\rho(X_{k})$
$\leq\sup_{y\in X_{k}}\lvert u(y)-u(x)\rvert$
$\leq Ch^{(\ell)},$
for some $C>0$, where in the last step we use that the function $u$ is Lipschitz continuous in $X_{k}$.
5 Numerical experiments
In this section we verify the performance of the refinement-based Christoffel sampling (RCS) algorithm numerically. Its implementation in Matlab and all code needed to reproduce the numerical experiments are made publicly available [22]. We use Matlab’s built-in function slicesample to sample from $\mu_{u}$, and we use either a simple Monte Carlo estimate or Matlab’s built-in function integral to estimate $\|u\|_{L_{\rho}^{1}(X)}$.
Some changes were made in the practical implementation compared to Algorithm 1, which favor the computational complexity while the empirical performance remains satisfactory. First, we discard $\log$-terms. More specifically, in each iterations we sample $C_{2}n^{\epsilon}$ points and construct the matrix $A$ via $(A)_{kj}=\overline{\phi_{j}(x_{k})}/\sqrt{C_{1}u(x_{k})}$. We choose $C_{1}=5$ and $C_{2}=5C_{1}$. At the end of the algorithm, we sample $C_{3}\|u\|_{L^{1}_{\rho}(X)}$ points with $C_{3}=10$. Second, we only take the first and the last two iterations into account for the update of $u$, i.e.,
$u_{(i)}(x)=\min\Bigg(\|k_{n}^{\epsilon}\|_{L^{\infty}(X)},\;\left\lVert\left(R_{(i-2)}^{*}\right)^{-1}\phi(x)\right\rVert_{2}^{2},\;\left\lVert\left(R_{(i-1)}^{*}\right)^{-1}\phi(x)\right\rVert_{2}^{2}\Bigg).$
With these changes, the computational complexity reduces to $\mathcal{O}\left(n^{3}\log\left(\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}\right)\right)$. For more details on the implementation, we refer to [22].
  1. A.
    Polynomials + weighted polynomials. We consider
    (21) $\Phi_{n}=\{p_{i}(x)\}_{i=0}^{n_{1}-1}\cup\{w(x)p_{i}(x)\}_{i=0}^{n_{2}-1},$
    where $n=n_{1}+n_{2}$ and $p_{i}$ is a polynomial basis function of degree $i$. Such bases are commonly used to approximate functions with known singular or oscillatory behaviour, which is incorporated in the weight function $w(x)$, see [4, Example 3] and references therein. The output of the RCS algorithm for $n=40$, X=[1,1]X=[-1,1] and $w(x)=\sqrt{x+1}$ is shown in Figure 2. In Figure 3, we consider the convergence of the least squares approximation of
    (22) $f(x)=\frac{\sqrt{x+1}}{1+5x^{2}}+\cos(5x)\qquad\text{for }x\in[-1,1]$
    in the span of $\Phi_{n}$ with the same weight function.
    (a) linear scale
    (b) logarithmic scale
    Figure 2: Illustration of the RCS algorithm for the polynomial + weighted polynomial basis (21) with $w(x)=\sqrt{x+1}$ and $n_{1}=n_{2}=20$ ($n=40$). The function $u$ computed by the RCS algorithm is compared to an approximation of $k_{n}^{\epsilon}$ obtained via the dense grid method.
    Figure 3: Convergence of the discrete least squares approximation to $f$ (22) in the polynomial + weighted polynomial basis (21) with $w(x)=\sqrt{x+1}$. A comparison is made between approximation based on uniformly random samples and samples outputted by the RCS algorithm. The line indicates the geometric mean and the shaded area shows the variance over ten repetitions.
  2. B.
    Rational approximation with preassigned poles. Recently, the lightning approximation scheme became popular for approximating functions with singularities at known locations, a setting which often occurs when solving PDEs on domains with corners [20]. As a model problem, consider approximation on $[0,1]$ of functions with an endpoint singularity at $x=0$ using the basis
    (23) $\Phi_{n}=\left\{\frac{-q_{i}}{x-q_{i}}\right\}_{i=1}^{n_{1}}\cup\{p_{i}(x)\}_{i=0}^{n_{2}-1}\qquad\text{with }q_{i}=-\exp\left(4\left(\sqrt{i}-\sqrt{n_{1}}\right)\right),$
    where $p_{i}$ is again a polynomial basis function of degree $i$, $n=n_{1}+n_{2}$ and $n_{2}=\mathcal{O}(\sqrt{n_{1}})$ [24]. Hence, we combine partial fractions with simple poles $q_{i}$, which cluster towards zero along the negative real axis in a tapered fashion [32], with a polynomial basis.
    Figure 5 shows the output of the RCS algorithm for different values of $n$. It becomes apparent that $k_{n}^{\epsilon}$ grows exponentially with the number of basis functions near the singularity. On Figure 5, the convergence behaviour of the discrete least squares fit to $f(x)=\sqrt{x}$ using $n_{2}=2\sqrt{n_{1}}$ is shown.
    Figure 4: Illustration of the RCS algorithm for the lightning + polynomial basis (23) for varying $n_{1}$ and $n_{2}$. The function $u$ computed by the RCS algorithm is compared to an approximation of $k_{n}^{\epsilon}$ obtained via the dense grid method.
    Figure 5: Convergence of the discrete least squares approximation to $f(x)=\sqrt{x}$ in the lightning + polynomial basis (23). A comparison is made between approximation based on uniformly random samples and samples outputted by the RCS algorithm. The line indicates the geometric mean and the shaded area shows the variance over ten repetitions.
    The lightning approximation scheme is extended to multiple dimensions in [8]. In order to illustrate the RCS algorithm in two dimensions, we consider the setting in [8, §4.2]. Namely,
    (24) $\Phi_{n}=\bigcup_{j=1}^{n_{1}}\left\{\frac{q_{j}p_{k}(x)p_{l}(y)}{Q(x,y)-q_{j}}\right\}_{(k,l)=(0,0)}^{(n_{2}-1,n_{2}-1)}\cup\{p_{k}(x)p_{l}(x)\}_{(k,l)=(0,0)}^{(n_{3}-1,n_{3}-1)}$
    on $X=[-2,2]^{2}$, where $p_{i}$ is again a polynomial basis function of degree $i$,
    $Q(x,y)=x^{3}-2x+1-y^{2}\qquad\text{and}\qquad q_{j}=\pm i\exp(-4(\sqrt{n_{1}}-\sqrt{j})),\;1\leq j\leq n_{1}.$
    Hence, we use $n=2n_{1}n_{2}^{2}+n_{3}^{2}$ basis functions. This basis is tailored to approximate functions with a curve of singularities along the elliptic curve $Q(x,y)=0$.
    In Figure 6, the output of the RCS algorithm is shown for $n_{1}=15$, $n_{2}=3$ and $n_{1}=10$ ($n=370$). The function $u$ is peaked along the elliptic curve, resulting in many clustered sample points close to the curve.
    (a) $u$
    (b) sample points
    Figure 6: Output of the RCS algorithm for the 2D lightning + polynomial basis (24) for $n_{1}=15,n_{2}=3$ and $n_{1}=10$ ($n=370$).
  3. C.
    Fourier extension of a surface. Non-orthogonal bases often arise when approximating on irregular domains. In such cases, many approximation methods first construct an orthonormal basis on a regular domain. However, when this basis is restricted to the irregular domain, it generally loses orthogonality. In the case of a Fourier basis, this approach is referred to as Fourier extension [28].
    We consider the extreme case of using a 3D Fourier basis to approximate a function on the surface of a box. More specifically, we use
    (25) $\Phi_{n}=\{\exp(2\pi i(k_{x}x+k_{y}y+k_{z}z))\}_{(k_{x},k_{y},k_{z})=(-n_{x},-n_{y},-n_{z})}^{(n_{x},n_{y},n_{z})}$
    to approximate functions on the surface of a cube inside the torus $1/2,1/2)^{3}$. The output of the RCS algorithm is shown in Figure 7. We find that the function $u$ computed by the RCS algorithm peaks near the edges and corners of the cube.
    (a) $u$
    (b) sample points
    Figure 7: Output of the RCS algorithm for the Fourier basis (25) restricted to the surface of a cube for $n_{x}=n_{y}=n_{z}=4$ ($n=729$).
  4. D.
    Extreme learning machines and neural networks. We consider a single-hidden-layer-feedforward neural network with fixed, randomly initialized hidden parameters. Such networks are often referred to as extreme learning machines (ELMs) [13]. When considering additive nodes, the basis equals
    (26) $\Phi_{n}=\{g(a_{i}x+b_{i})\}_{i=1}^{n}\qquad$
    where $a_{i}\in\mathbb{R}^{d},b_{i}\in\mathbb{R}$ are randomly intialized and $g$ is a bounded nonconstant piecewise continous activation function. Universal approximation properties for such networks on compact input sets are proven in [26].
    We consider (26) on $0,1]^{2}$ with $a_{i}$ uniformly randomly sampled in $1,1]^{2}$ and $b_{i}$ uniformly randomly sampled in $[0,1]$, $g$ sigmoidal $\left(g(x)=\frac{1}{1+\exp(-x)}\right)$ and $n=600$. Figure 8 depicts the first two basis function associated with this ELM and the output of the RCS algorithm.
    Note that similar techniques can be applied even when the hidden nodes are not fixed. Specifically, the output of a deep neural network can be viewed as a linear combination of the outputs of basis functions associated with the nodes in the penultimate layer. This perspective, often called the adaptive basis viewpoint, naturally leads to hybrid training procedures: the weights in the last layer are learned via least squares, while all other weights are optimized using gradient descent [12]. In [3], the least squares component is further enhanced using Christoffel sampling.
    (a) $\phi_{1}$
    (b) $\phi_{2}$
    (c) $u$
    (d) sample points
    Figure 8: First basis functions and output of the RCS algorithm for one instance of an ELM (26) on $X=[0,1]^{2}$, with sigmoidal activation function, $n=600$, $a_{i}$ uniformly randomly sampled in $1,1]^{2}$ and $b_{i}$ uniformly randomly sampled in $[0,1]$.
Appendix A Discrete result
In this appendix, we prove a discrete result very similar to [10, Theorem 2] (for leverage scores) and [11, Lemma 6.2] (for ridge leverage scores with variable regularization parameter). The main difference is that the following result holds for ridge leverage scores with a fixed regularization parameter, which does not pose any new challenges compared to known results. We will use the following notation:
$\tau_{k}^{\epsilon}(A)$ $=a_{k}^{\top}\left(A^{*}A+\epsilon^{2}I\right)^{-1}(a_{k}^{\top})^{*}$
$n^{\epsilon}(A)$ $=\sum_{k=1}^{m}\tau_{k}^{\epsilon}(A)=\sum_{i=1}^{n}\frac{\lambda_{i}(A^{*}A)}{\lambda_{i}(A^{*}A)+\epsilon^{2}}$
where $a_{k}^{\top}$ is the $k$-th row of the matrix $A$. This Lemma is used to prove Lemma 4.3.
Lemma A.1.
Let $A\in\mathbb{C}^{m\times n}$, $u\in\mathbb{R}^{m}$ with $u_{k}>0$ for all $1\leq k\leq m$ and $\epsilon>0$, then there exists a diagonal matrix $W\in\mathbb{R}^{m\times m}$ with $0\preceq W\preceq I$ such that
(27) $\tau_{k}^{\epsilon}(WA)\leq u_{k},\qquad 1\leq k\leq m,$
and
(28) $\sum_{k:W_{kk}\neq 1}u_{k}\leq n^{\epsilon}(A).$
Proof A.2.
The proof is very similar to that of [10, Theorem 2]. Note that for $\epsilon>0$, every ridge leverage score is strictly smaller than one since $(WA)^{*}WA+\epsilon^{2}I\succ(WA)^{*}WA$ and, hence,
(29) $\tau_{k}^{\epsilon}(WA)<\tau_{k}(WA)\leq 1.$
The “whack-a-mole” algorithm can, therefore, be adapted to Algorithm 2. Similarly, for all $j\geq 0$, we let $W^{(j)}$ be the value of the $j$-th update to the weight and we show that $\overline{W}=\lim_{j\to\infty}W^{(j)}$ meets the conditions. Using the same arguments as in the proof of [10, Theorem 2], it follows from [10, Lemma 5] (which can easily be shown to hold also for $\tau_{k}^{\epsilon}$) that $W^{(j)}$ exists, its entries are nonnegative for all $j\geq 0$ and $\overline{W}$ exists. Furthermore, since $\epsilon>0$ and $W^{(j)}$ is nonnegative, it is easy to see that $\tau_{k}^{\epsilon}(WA)$ is continuous in $W$. Hence, by construction we have
$\tau_{k}^{\epsilon}(\overline{W}A)=\lim_{j\to\infty}\tau_{k}^{\epsilon}(W^{(j)}A)\leq u_{k},\qquad\forall k.$
It only remains to show that (28) holds for $W=\overline{W}$. Let $j$ be the first iteration such that $W_{kk}^{(j)}\neq 1$ for any $k$ such that $\overline{W}_{kk}\neq 1$. Following [10, Lemma 5], decreasing the weight of one row can only increase the ridge leverage scores of all other rows. This implies that
$\sum_{k:\overline{W}_{kk}\neq 1}u_{k}$ $\leq\sum_{k:\overline{W}_{kk}\neq 1}\tau_{k}^{\epsilon}(W^{(j)}A)$
$\leq n^{\epsilon}(W^{(j)}A)$
$=\sum_{i=1}^{n}\frac{\lambda_{i}((W^{(j)}A)^{*}(W^{(j)}A))}{\lambda_{i}((W^{(j)}A)^{*}(W^{(j)}A))+\epsilon^{2}}$
$\leq\sum_{i=1}^{n}\frac{\lambda_{i}(A^{*}A)}{\lambda_{i}(A^{*}A)+\epsilon^{2}}$
$=n^{\epsilon}(A),$
where in the penultimate step we use that $(W^{(j)}A)^{*}W^{(j)}A\preceq A^{*}A$ and, hence, $\lambda_{i}((W^{(j)}A)^{*}W^{(j)}A)\leq\lambda_{i}(A^{*}A),\forall i$ [25, §9-5], as well as the fact that $\lambda/(\lambda+\epsilon^{2})$ is monotonically increasing for nonnegative $\lambda$.
Algorithm 2 Whack-a-mole algorithm
1:Initialize $W=I$
2:while true do
3:  for $k=1$ to $m$ do
4:   if $\tau_{k}^{\epsilon}(WA)\geq u_{k}$ then
5:      Decrease $W_{kk}$ so that $\tau_{k}^{\epsilon}(WA)=u_{k}$.
6:   end if
7:  end for
8:end while
9:return $W$
Appendix B Auxiliary results
Lemma B.1.
Let $X\subseteq\mathbb{R}^{d}$, $\{\phi_{i}\}_{i=1}^{n}\subset L^{2}_{\rho}(X)$, $\epsilon>0$ and $v:X\to\mathbb{R}$ be a function satisfying
$0\leq v(x)\leq 1,\qquad\forall x\in X.$
Then,
$n^{\epsilon}(v\Phi)\leq n^{\epsilon}(\Phi).$
Proof B.2.
Denote the gram matrices with respect to the $L^{2}_{\rho}(X)$ inner product associated with the sets $\Phi$ and $v\Phi$ by $G$ and $G^{v}$, respectively. Since
$(G)_{ij}-(G^{v})_{ij}=\int_{X}(1-v(x)^{2})\phi_{i}(x)\phi_{j}(x)d\rho=\langle\sqrt{1-v^{2}}\phi_{i},\sqrt{1-v^{2}}\phi_{j}\rangle_{L^{2}_{\rho}(X)}$
one has $G-G^{v}\succeq 0\Rightarrow G\succeq G^{v}$. This implies that $\lambda_{i}(G^{v})\leq\lambda_{i}(G),\forall i$ [25, §9-5], which allows us to bound
$n^{\epsilon}(v\Phi)=\sum_{i=1}^{n}\frac{\lambda_{i}(G^{v})}{\lambda_{i}(G^{v})+\epsilon^{2}}\leq\sum_{i=1}^{n}\frac{\lambda_{i}(G)}{\lambda_{i}(G)+\epsilon^{2}}=n^{\epsilon}(\Phi),$
using the fact that $\lambda/(\lambda+\epsilon^{2})$ is monotonically increasing for nonnegative $\lambda$.
Lemma B.3.
Let $X\subseteq\mathbb{R}^{d}$, $\{\phi_{i}\}_{i=1}^{n}\subset L^{2}_{\rho}(X)$, $M\in\mathbb{C}^{n\times n}$ positive semidefinite, $\epsilon>0$ and $\beta\geq 1$. Then,
$k^{\epsilon}(x;\Phi,M)\leq\beta k^{\epsilon}(x;\Phi,\beta M).$
Proof B.4.
One has
$k^{\epsilon}(x;v\Phi,M)=\phi(x)^{*}(G+\epsilon^{2}I)^{-1}\phi(x)=\beta\phi(x)^{*}(\beta G+\beta\epsilon^{2}I)^{-1}\phi(x).$
Furthermore, $\beta G+\beta\epsilon^{2}I\succeq\beta G+\epsilon^{2}I$ since $\beta\geq 1$ and, hence,
$k^{\epsilon}(x;v\Phi,M)\leq\beta\phi(x)^{*}(\beta G+\epsilon^{2}I)^{-1}\phi(x)=\beta k^{\epsilon}(x;\Phi,\beta M)$
as required.
Acknowledgments
The authors wish to thank Daan Huybrechs and Christopher Musco for fruitful discussions on the topic.